OpenStudy (anonymous):

For spring break this year your family decided to visit Michigan, but it snows the first weekend you are there. You decide to make a snowman. If the bottom ball is 3 feet across, the middle ball is 2 feet across, and the top is 1 foot across, how much snow is needed to build the whole snowman? Round your answer to the nearest cubic foot. A) 19 cubic feet of snow B) 31 cubic feet of snow C) 511 cubic feet of snow D) 1022 cubic feet of snow

OpenStudy (anonymous):

\[V=\frac{ 4 }{ 3 }{\pi}r^3\]Add each of the volumes together

OpenStudy (anonymous):

A) 19 cubic feet equation: (1ft x3.14)+(2ft x3.14)+(3ft x3.14)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I stand corrected he is correct but used the wrong formula. I would like an explanation how you did it. or did you get lucky

OpenStudy (anonymous):

@ChmE lol which equation are you suppose to use

OpenStudy (anonymous):

The one I posted above. You have to find the volume of each sphere and add them together. Not just add 1pi + 2pi + 3pi. The answer you get is 18.85 my way. That round to 19

OpenStudy (anonymous):

the way you explained it is way easier

OpenStudy (anonymous):

OMG. \[1\pi + 2\pi + 3\pi \]Gives the exact same answer as my way. But that is a very unconventional way to do it. pi does not equal (4/3)pi(.5^2) but when you add it all up it yields the same answer. This appears to be a shortcut to adding volumes of spheres with this unique pattern. I'd be curious to know if it works for all circumstances. I'd stick with the correct volume formula to be safe, but this was interesting to see for me.

OpenStudy (anonymous):

Thanks :)